# Personalized Subgraph Federated Learning with Differentiable Auxiliary Projections

### v1

@@ C.3.1 Local Training Analysis
Then we can insert Eq. 35 and Eq. 36 into Eq. 34 and take expectation as follows:
@@
Because $\eta\leq 1/2\mathscr{L}$ we have $21-\mathscr{L}\eta/2\geq 1/$ 2, hence:
@@
Based on the assumption that each local objective $\mathcal{L}_{k}$ satisfies the $\mu$-PL  [^2] Here, we slightly abuse the notation $\mu$, which was previously introduced in Sec. C.1.1 with a different meaning. (Polyak-Lojasiewicz) condition (Polyak, 1963) iff
@@
We define the gap as $\Delta^{q-1}_{k}:=\mathcal{L}_{k}(\Psi_{k}^{q-1})-\mathcal{L}^{\star}_{k}$. Taking the total expectation and iterating Eq. 40 $Q$ times, we have:
@@ C.3.2 Effect of Global Kernel-Based Aggregation
Define $\Psi^{\mathrm{loc},(t-1)}_{k}:=\Psi^{Q}_{k}$ which means the local client parameters after $Q$ local training iterations. Let $\mathbf{f}^{(t-1)}:=\left[f_{1}^{(t-1)},\cdots,f_{K}^{(t-1)}\right]^{\top}$, where $f_{k}^{(t-1)}=\mathcal{L}_{k}(\Psi_{k}^{\mathrm{loc},(t-1)})-\mathcal{L}^{\star}_{k}$. Recall our proposed personalized aggregation scheme in Eq. 8, it can be rewritten as:
@@
Since each new parameter is a convex combination Eq. 44, based on the Jensen’s inequality and $\mathscr{L}$-smoothness assumption in Eq. 33, the following inequality holds:
@@
Let $\mathbf{p}=[p_{1},\dots,p_{K}]^{\top}$, Eq. 45 subtracts $\mathcal{L}_{k}^{\star}$ and multiply by $p_{k}$, and sum over $k$, we can reach:
@@
Let $c:=1/\sqrt{\min_{k}p_{k}}\leq\sqrt{K}$, we can combine Eq. 50 and Eq. 51 to get the following inequality:
@@
Then by squaring both sides of Eq. 52 and use $\sqrt{\bar{f}}\leq 1+\bar{f}$, we have:
@@
where the last inequality employs $\bar{f}^{(t-1)}\leq(1-\rho)^{-1}\bar{f}^{(t-1)}$ which is trivial for $10<\rho<$ 1. Taking expectations, we can substitute Eq. 53 in to Eq. 48 to obtain:
@@
Given Eq. 43 and $\bar{f}^{(t-1)}=\sum_{k}p_{k}\mathbb{E}\left[f_{k}^{(t-1)}\right]$, we can bound $\mathbb{E}[f^{(t-1)}_{k}]$ by:
@@
By plugging Eq. 56 into Eq. 54, we have:
@@
By setting $\gamma:=(1-\eta\mu)^{Q}$ where $10<\gamma<$ 1, we can unroll Eq. 58 as:
@@
By inserting the above inequality into Eq. 59 and simplify, we can easily get:
@@
Recovering $\gamma^{T}=(1-\eta\mu)^{QT}$ gives exactly Eq. 12 in Theorem 3.3, which concludes the proof.
@@ D.2 Implementation Details
Following the standard FL settings, in our FedAux, both the local client models and the global server model adopt the same backbone architecture. We employ MaskedGCN (Baek et al., 2023) to generate node embeddings and sweep the number of GCN layers over $L\in\{1,2,3\}$. The hidden dimension is selected from $d^{\prime}\in\{64,128,256\}$, and dropout probabilities are set to 0.5. The auxiliary projection vector $\bm{a}$ is initialised from a Gaussian distribution in $\mathbb{R}^{d^{\prime}}$. The similarity–temperature parameter $\alpha$ is set to 10, and the bandwidth $\sigma$ is fixed to 1. For the FL schedule, we run $T=100$ communication rounds with $Q=1$ local epoch on the smaller citation datasets (Cora, Citeseer, PubMed). On all other datasets, we set the total number of rounds to $T=200$ and the number of local epochs per round to $Q=2$. All experiments are executed on a workstation equipped with an NVIDIA Tesla V100 SXM2 GPU (32 GB) running CUDA 12.4.
D.3 Quantify Non-IIDness in Federated Graph Datasets
@@
where $R_{k}=\frac{1}{2}(P_{k}+P)$ denotes the mid-point distribution between $P_{k}$ and $P$ with $R_{k}(y\in Y)=\frac{1}{2}\left[P_{k}(y)+P(y)\right]$. Pinsker’s inequality (Csiszár & Körner, 2011) gives that $\textsc{Jsd}\in[0,\log 2]$. A small Jsd indicates that each client’s label distribution closely matches the global prior, so the partition is effectively IID. As the Jsd increases, local class proportions deviate more sharply from the global mixture, making individual clients progressively class-specific and therefore increasingly subject to statistical non-IIDness.
@@
Label skew alone may underestimate heterogeneity when covariate shift is strong. To quantify covariate-shift–induced heterogeneity, we measure how far the embedding distribution of each client deviates from that of every other client in an embedding space that reflects graph structure and attributes. We first obtain node embeddings for all nodes with a simple neighbor aggregation $\mathbf{Z}_{k}=\mathbf{A}_{k}\mathbf{X}_{k}=\{z_{k,i}\}^{N_{k}}_{i=1}$ in each client. Let $\mathbf{Z}_{k}$ be the empirical distribution of embeddings held by client $G_{k}$. We define the graph-distribution discrepancy of a $K$-client partition as the mean pair-wise maximum mean discrepancy (MMD) as:
@@
where $\phi(\cdot)$ is the canonical feature map of a Gaussian RBF kernel with the bandwidth fixed with the median pairwise distance to ensure comparability across datasets. Eq. 63 evaluates to zero when all clients share an identical embedding distribution (IID) and increases monotonically with covariate divergence.
Based on the above two perspectives, we can quantify the degree of non-IIDness by summing Jsd and Mmd as $\xi=\textsc{Jsd}+\textsc{Mmd}$, where a higher $\xi$ indicates a higher degree of non-IIDness. We show the degree of non-IIDness of all datasets under different numbers of clients in Table 4.
@@ D.4 Synthetic Graph for Client Similarity Estimation
We first generate an SBM graph with 3000 nodes that are uniformly divided into $K=20$ equal-sized blocks as clients. Inter-client edges are added with probability $P^{\mathrm{inter}}=0.02$. To inject structural non-IIDness, the 20 clients are grouped into five super-clusters $\{\mathcal{G}_{i}\}_{i=1}^{5}$ (four clients per group). For every client belonging to $\mathcal{G}_{i}$, we draw intra-client edges with probability $P^{\mathrm{intra}}_{i}=0.15\times i$, so clients within the same group are structurally IID, while clients across groups are non-IID. To inject label and feature non-IIDness, we assign 5 labeled classes to all nodes. Nodes owned by the group $\mathcal{G}_{i}$ receive label $i-1$ with high probability 0.8, and the remaining 0.2 mass is distributed uniformly over the other four labels. Each node’s feature vector is the one-hot encoding of its label. In this way, clients in the same group are IID in both label and feature space, while clients from different groups exhibit pronounced distributional shifts. This controlled setting allows us to test whether the learned APVs can cluster clients that are genuinely similar. We directly compute the similarity between clients’ mean-pooling embeddings without considering privacy in Fig. 3(a) as the ground-truth similarity. Fig. 3(b)-3(d) are similarities computed by APVs, learnable weights of the readout layer, and functional embedding (Baek et al., 2023).
@@ E.1 Ablation Study
We conduct an ablation study in Table 5 validate our motivation and design, with the following two ablations: (i) replacing our proposed continuous aggregation scheme over the $\bm{a}_{k}$-space defined in Eq. 6 with the Conv1D operation applied on the hard-sorted embeddings as introduced by Liu et al. (2021), yielding a variant FedAuxhard, and (ii) removing our server-side APV-based personalized aggregation, instead using simple averaging to aggregate local models into a global model, leading to the baseline FedAvgmask. Note that the FedAvgmask variant differs from the standard FedAvg used in Table 1, where FedAvgmask adopts MaskedGCN (Baek et al., 2023) as the GNN backbone, whereas the standard FedAvg utilizes a conventional GCN architecture. Compared with FedAuxhard, our FedAux consistently obtains higher accuracy. This empirical gain confirms our claim that although the hard-sorting scheme proposed by Liu et al. (2021) does allow gradient flow to optimize the APV, the underlying discrete permutation remains non-differentiable and therefore restricts the capacity of the APV to adapt. By using our proposed continuous aggregation with a fully differentiable kernel operator, FedAux enables smoother gradient flow, allowing the APV and the local GNN to co-evolve optimally. Besides, FedAux and FedAvgmask both adopt MaskedGCN as backbone, while FedAux personalized aggregates local models based on the APV similarity, while FedAvgmask aggregates local models to a single global model. Results show that FedAux outperforms FedAvgmask on all datasets, which demonstrates that exploiting APV-based personalized aggregation allows the federation to respect non-IID data distributions and learn more effective client-specific models.
@@ E.2 Hyperparameter Analysis
In our model, the number of GCN layers is selected from $L\in\{1,2,3\}$, and the dimension of the hidden layers is selected from $d^{\prime}\in\{64,128,256\}$. In Fig. 6, we show all the hyperparameter combinations on the Cora dataset for different client counts. It is evident that FedAux consistently achieves the highest performance with $L=2$ across all federated settings. For the hidden dimension, when the number of clients is not large, FedAux requires a relatively small hidden dimension, while with 20 clients, a hidden dimension of 256 yields the best results.
@@
In Eq. 7, the similarity temperature parameter $\alpha$ is introduced to modulate the sharpness of the similarity distribution. To evaluate the sensitivity of FedAux to this hyperparameter, we test $\alpha\in\{0.1,1,10\}$ across varying numbers of clients and report the resulting accuracy in Fig. 7. The results show that while some configurations achieve optimal accuracy at different values of $\alpha$, for example, $\alpha=0.1$ is optimal for 20 clients on Cora and $\alpha=10$ is optimal for 20 clients on Pubmed, directly setting $\alpha=1$ consistently provides satisfactory performance across different datasets and client counts.

### v2

@@ C.3.1 Local Training Analysis
Then we can insert Eq.˜35 and Eq.˜36 into Eq.˜34 and take expectation as follows:
@@
Because $\eta\leq 1/2\mathscr{L}$ we have $1-\mathscr{L}\eta/2\geq 1/2$, hence:
@@
Based on the assumption that each local objective $\mathcal{L}_{k}$ satisfies the $\mu$-PL  [^2] Here, we slightly abuse the notation $\mu$, which was previously introduced in Section C.1.1 with a different meaning. (Polyak-Lojasiewicz) condition [23] iff
@@
We define the gap as $\Delta^{q-1}_{k}:=\mathcal{L}_{k}(\Psi_{k}^{q-1})-\mathcal{L}^{\star}_{k}$. Taking the total expectation and iterating Eq.˜40 $Q$ times, we have:
@@ C.3.2 Effect of Global Kernel-Based Aggregation
Define $\Psi^{\mathrm{loc},(t-1)}_{k}:=\Psi^{Q}_{k}$ which means the local client parameters after $Q$ local training iterations. Let $\mathbf{f}^{(t-1)}:=\left[f_{1}^{(t-1)},\cdots,f_{K}^{(t-1)}\right]^{\top}$, where $f_{k}^{(t-1)}=\mathcal{L}_{k}(\Psi_{k}^{\mathrm{loc},(t-1)})-\mathcal{L}^{\star}_{k}$. Recall our proposed personalized aggregation scheme in Eq.˜8, it can be rewritten as:
@@
Since each new parameter is a convex combination Eq.˜44, based on the Jensen’s inequality and $\mathscr{L}$-smoothness assumption in Eq.˜33, the following inequality holds:
@@
Let $\mathbf{p}=[p_{1},\dots,p_{K}]^{\top}$, Eq.˜45 subtracts $\mathcal{L}_{k}^{\star}$ and multiply by $p_{k}$, and sum over $k$, we can reach:
@@
Let $c:=1/\sqrt{\min_{k}p_{k}}\leq\sqrt{K}$, we can combine Eq.˜50 and Eq.˜51 to get the following inequality:
@@
Then by squaring both sides of Eq.˜52 and use $\sqrt{\bar{f}}\leq 1+\bar{f}$, we have:
@@
where the last inequality employs $\bar{f}^{(t-1)}\leq(1-\rho)^{-1}\bar{f}^{(t-1)}$ which is trivial for $0<\rho<1$. Taking expectations, we can substitute Eq.˜53 in to Eq.˜48 to obtain:
@@
Given Eq.˜43 and $\bar{f}^{(t-1)}=\sum_{k}p_{k}\mathbb{E}\left[f_{k}^{(t-1)}\right]$, we can bound $\mathbb{E}[f^{(t-1)}_{k}$] by:
@@
By plugging Eq.˜56 into Eq.˜54, we have:
@@
By setting $\gamma:=(1-\eta\mu)^{Q}$ where $0<\gamma<1$, we can unroll Eq.˜58 as:
@@
By inserting the above inequality into Eq.˜59 and simplify, we can easily get:
@@
Recovering $\gamma^{T}=(1-\eta\mu)^{QT}$ gives exactly Eq.˜12 in Theorem˜3.3, which concludes the proof.
@@ D.2 Implementation Details
Following the standard FL settings, in our FedAux, both the local client models and the global server model adopt the same backbone architecture. We employ MaskedGCN [3] to generate node embeddings and sweep the number of GCN layers over $L\in\{1,2,3\}$. The hidden dimension is selected from $d^{\prime}\in\{64,128,256\}$, and dropout probabilities are set to 0.5. The auxiliary projection vector $\bm{a}$ is initialised from a Gaussian distribution in $\mathbb{R}^{d^{\prime}}$. The similarity–temperature parameter $\alpha$ is set to 10, and the bandwidth $\sigma$ is fixed to 1. For the FL schedule, we run $T=100$ communication rounds with $Q=1$ local epoch on the smaller citation datasets (Cora, Citeseer, PubMed). On all other datasets, we set the total number of rounds to $T=200$ and the number of local epochs per round to $Q=2$. All experiments are executed on a workstation equipped with an NVIDIA Tesla V100 SXM2 GPU (32 GB) running CUDA 12.4.
D.3 Quantifying Non-IIDness in Federated Graph Datasets
@@
where $R_{k}=\frac{1}{2}(P_{k}+P)$ denotes the mid-point distribution between $P_{k}$ and $P$ with $R_{k}(y\in Y)=\frac{1}{2}\left[P_{k}(y)+P(y)\right]$. Pinsker’s inequality [6] gives that $\textsc{Jsd}\in[0,\log 2]$. A small Jsd indicates that each client’s label distribution closely matches the global prior, so the partition is effectively IID. As the Jsd increases, local class proportions deviate more sharply from the global mixture, making individual clients progressively class-specific and therefore increasingly subject to statistical non-IIDness.
@@
Label skew alone may underestimate heterogeneity when covariate shift is strong. To quantify covariate‐shift–induced heterogeneity, we measure how far the embedding distribution of each client deviates from that of every other client in an embedding space that reflects graph structure and attributes. We first obtain node embeddings for all nodes with a simple neighbor aggregation $\mathbf{Z}_{k}=\mathbf{A}_{k}\mathbf{X}_{k}=\{z_{k,i}\}^{N_{k}}_{i=1}$ in each client. Let $\mathbf{Z}_{k}$ be the empirical distribution of embeddings held by client $G_{k}$. We define the graph-distribution discrepancy of a $K$-client partition as the mean pair-wise maximum mean discrepancy (MMD) as:
@@
where $\phi(\cdot)$ is the canonical feature map of a Gaussian RBF kernel with the bandwidth fixed with the median pairwise distance to ensure comparability across datasets. Eq.˜63 evaluates to zero when all clients share an identical embedding distribution (IID) and increases monotonically with covariate divergence.
Based on the above two perspectives, we can quantify the degree of non-IIDness by summing Jsd and Mmd as $\xi=\textsc{Jsd}+\textsc{Mmd}$, where a higher $\xi$ indicates a higher degree of non-IIDness. We show the degree of non-IIDness of all datasets under different numbers of clients in Table˜5.
@@ D.4 Synthetic Graph for Client Similarity Estimation
We first generate an SBM graph with 3000 nodes that are uniformly divided into $K=20$ equal-sized blocks as clients. Inter-client edges are added with probability $P^{\mathrm{inter}}=0.02$. To inject structural non-IIDness, the 20 clients are grouped into five super-clusters $\{\mathcal{G}_{i}\}_{i=1}^{5}$ (four clients per group). For every client belonging to $\mathcal{G}_{i}$, we draw intra-client edges with probability $P^{\mathrm{intra}}_{i}=0.15\times i$, so clients within the same group are structurally IID, while clients across groups are non-IID. To inject label and feature non-IIDness, we assign 5 labeled classes to all nodes. Nodes owned by the group $\mathcal{G}_{i}$ receive label i1i-1 with high probability 0.8, and the remaining 0.2 mass is distributed uniformly over the other four labels. Each node’s feature vector is the one-hot encoding of its label. In this way, clients in the same group are IID in both label and feature space, while clients from different groups exhibit pronounced distributional shifts. This controlled setting allows us to test whether the learned APVs can cluster clients that are genuinely similar. We directly compute the similarity between clients’ mean-pooling embeddings without considering privacy in Fig.˜3(a) as the ground-truth similarity. Fig.˜3(b)-3(d) are similarities computed by APVs, learnable weights of the readout layer, and functional embedding [3].
@@ E.1 Ablation Study
We conduct an ablation study in Table˜6 validate our motivation and design, with the following two ablations: (i) replacing our proposed continuous aggregation scheme over the $\bm{a}_{k}$-space defined in Eq.˜6 with the $\mathrm{Conv1D}$ operation applied on the hard-sorted embeddings as introduced by [18], yielding a variant FedAuxhard, and (ii) removing our server-side APV-based personalized aggregation, instead using simple averaging to aggregate local models into a global model, leading to the baseline FedAvgmask. Note that the FedAvgmask variant differs from the standard FedAvg used in Table˜1, where FedAvgmask adopts MaskedGCN [3] as the GNN backbone, whereas the standard FedAvg utilizes a conventional GCN architecture. Compared with FedAuxhard, our FedAux consistently obtains higher accuracy. This empirical gain confirms our claim that although the hard-sorting scheme proposed by [18] does allow gradient flow to optimize the APV, the underlying discrete permutation remains non-differentiable and therefore restricts the capacity of the APV to adapt. By using our proposed continuous aggregation with a fully differentiable kernel operator, FedAux enables smoother gradient flow, allowing the APV and the local GNN to co-evolve optimally. Besides, FedAux and FedAvgmask both adopt MaskedGCN as backbone, while FedAux personalized aggregates local models based on the APV similarity, while FedAvgmask aggregates local models to a single global model. Results show that FedAux outperforms FedAvgmask on all datasets, which demonstrates that exploiting APV-based personalized aggregation allows the federation to respect non-IID data distributions and learn more effective client-specific models.
@@ E.2 Hyperparameter Analysis
In our model, the number of GCN layers is selected from $L\in\{1,2,3\}$, and the dimension of the hidden layers is selected from $d^{\prime}\in\{64,128,256\}$. In Fig.˜7, we show all the hyperparameter combinations on the Cora dataset for different client counts. It is evident that FedAux consistently achieves the highest performance with $L=2$ across all federated settings. For the hidden dimension, when the number of clients is not large, FedAux requires a relatively small hidden dimension, while with 20 clients, a hidden dimension of 256 yields the best results.
@@
In Eq.˜7, the similarity temperature parameter $\alpha$ is introduced to modulate the sharpness of the similarity distribution. To evaluate the sensitivity of FedAux to this hyperparameter, we test $\alpha\in\{0.1,1,10\}$ across varying numbers of clients and report the resulting accuracy in Fig.˜8. The results show that while some configurations achieve optimal accuracy at different values of $\alpha$, for example, $\alpha=0.1$ is optimal for 20 clients on Cora and $\alpha=10$ is optimal for 20 clients on Pubmed, directly setting $\alpha=1$ consistently provides satisfactory performance across different datasets and client counts.
Appendix F Limitations and Future Work
The main limitation of our work is that while the kernel‑based continuous aggregation scheme lets each client learn a fully differentiable, data‑driven ordering on the APV, it requires evaluating the pairwise kernel for every node pair in the local client subgraph. Thus, the per‑client computational cost scales quadratically, $\mathcal{O}(N_{k}^{2})$, which is acceptable for small and medium subgraphs but can dominate run‑time on very large clients. Our future work focuses on addressing this scalability bottleneck. One promising direction is to approximate the dense kernel with low‑rank factorizations, such as Nyström approximation, so that complexity scales linearly in $N_{k}$ with controllable error. We stress that our core contribution is a privacy‑preserving mechanism for personalized clustering clients via the APV, and reducing local computational load is a complementary line of future research.